# Metaverse Identity: Core Principles and Critical Challenges

### v1

@@ Abstract
This paper explores the core principles that should guide the construction and governance of identity in the metaverse and identifies the critical challenges that need to be addressed. Drawing on multidisciplinary theories and perspectives, we propose two core principles for metaverse identity: Equivalence and Alignment, and Fusion and Expansiveness. The first principle contends that metaverse identities should be consistent with real-world identities in terms of norms and standards, which is crucial for establishing guidelines and safeguarding rights. The second principle emphasizes the necessity for seamless integration and boundless expansion of metaverse identities, transcending real-world limitations to accommodate diverse needs and foster inclusive participation. We argue that these two principles are vital for ensuring the accountability, inclusiveness, and consistency of identity in the metaverse. We also identify five critical challenges: Identity Interoperability, Legal Implications, Privacy and Identity Management, Deepfakes and Synthetic Identities, and Identity Fragmentation and Psychological Well-being. We discuss potential strategies to navigate these challenges. The paper concludes by underscoring the importance of a proactive and collaborative approach to shaping the future of metaverse identity. As the metaverse continues to evolve, it is imperative that we cultivate a thorough understanding of the principles and challenges surrounding identity in this uncharted territory and work collectively to build a metaverse that fosters responsible identity construction and expression.
@@
Metaverse, Identity, Governance, Principles, Interoperability, Privacy, Deepfakes, Synthetic Identity, Identity Fragmentation
journal: ArXiv
@@
[Liang Yang]organization=Division of Emerging Interdisciplinary Areas, The Hong Kong University of Science and Technology,country=Hong Kong SAR
[…]
[Yan Xu]organization=Information Systems, Business Statistics and Operations Management, The Hong Kong University of Science and Technology,country=Hong Kong SAR
[…]
[Pan Hui]organization=Computational Media and Arts Thrust, The Hong Kong University of Science and Technology (Guangzhou),city=Guangzhou, country=China
@@
This study articulates two core principles for the governance of identity in the metaverse: "Equivalence and Consistency" and "Integration and Expandability."
Furthermore, the paper systematically identifies and examines five critical challenges that impede the effective management of identity in the metaverse.
@@ 1 Introduction
The metaverse is envisioned as the future of the Internet, aspiring to offer a spatial and social Internet experience that seamlessly blend the physical and digital realms using both existing and emerging technologies (Venugopal et al.,, 2023; Hollensen et al.,, 2022; Koohang et al.,, 2023). Over the past few decades, and particularly in the last three years, scholarly research on the metaverse has extensively covered various dimensions, including technological advancements (Park and Kim, 2022b, ; Xu et al.,, 2022; Huynh-The et al.,, 2022; Fu et al.,, 2023), ecosystem dynamics (Lee et al.,, 2021), social impacts (Hennig-Thurau et al.,, 2022; Šímová et al.,, 2023; Park and Kim, 2022b, ), economic potential (Bhattacharya et al.,, 2023), and user behavior and experiences (Riva and Wiederhold,, 2022; Park and Kim, 2022a, ; Buhalis et al.,, 2022).
@@
However, the development and adoption of the metaverse also presents new challenges. In January, the International Criminal Police Organization(INTERPOL) released a research report titled "Metaverse - A Law Enforcement Perspective." This report underscores that, despite its nascent stages, the metaverse already poses urgent technological, social, and ethical challenges that demands immediate attention. Notably, the metaverse has facilitated the emergence "meta-crimes," new types of criminal activities that are increasingly concerning as immersive worlds become more integral to our daily lives. Moreover, the report highlight significant governance and policy issues, with the construction of identity within the metaverse being a the primary one. It notes that the anonymity of identities within the metaverse has been exploited for illegal activities such as trafficking stolen goods or promoting unlawful behavior. Additionally, the legal status of avatars remains undefined, prompting the international community to actively explore their legal responsibilities and definitions (INTERPOL,, 2024).
@@
In March, the World Economic Forum(WEF) released a report entitled "Metaverse Identity: Defining the Self in a Blended Reality." This report elucidates that the metaverse acts as a conduit merging the digital and physical worlds, thereby revolutionizing interactions with information, individuals, and environments. Within this transformative context, identity stands out as a critical element, essential for promoting inclusivity, fairness, accessibility, security, and privacy within the metaverse. The report warns that neglecting the evolution of identity in the metaverse could result in the perpetuation of existing internet flaws. Poorly designed metaverse identities could adversely affect social mobility in the physical world—amidst our increasing reliance on digital platforms—and hinder the secure and private identification, authentication, and verification of individuals or digital entities. Furthermore, these identities may unintentionally endow technology with human-like attributes, cultivating excessive trust and potentially leading to various psychological and emotional harms (WEF,, 2024).
@@
This paper aims to systematically review academic research concerning Metaverse Identity, to explore its intrinsic characteristics and impacts, and to propose core principles that aid stakeholders in intuitively understanding both Metaverse Identity and its broader impacts. Additionally, this study seeks to identify the current critical challenges, aiming to provide theoretical references and insights for technological innovations, ethical standards, and policy developments related to Metaverse Identity. Employing a comprehensive, primarily qualitative methodology, this study will integrate literature reviews, textual analysis, comparative studies, and case studies to construct a multidimensional analytical framework. This diversified methodological approach, proven effective for complex subjects, aligns with strategies used by other scholars in studies of virtual identities (Rani,, 2022; Mitrushchenkova,, 2023). Acknowledging the inherent limitations due to the expansive scope of our analysis, this paper seeks to catalyze further interdisciplinary dialogue and research. It is envisioned that future studies will build on this groundwork to provide more profound analyses and robust solutions.
@@ 2 Literature Review
The investigation of Identity represents a crucial philosophical inquiry into human existence. At a fundamental level, Identity encompasses personal details about who they are such as name, age, gender, place of origin, and nationality. However, the Cambridge Dictionary expands this definition to include it as "the fact of being, or feeling that you are, a particular type of person, organization, etc.; the qualities that make a person, organization, etc. different from others" [^1] https://dictionary.cambridge.org/dictionary/english/identity. This broader perspective aligns with philosophical explorations of the ’Self’, epitomized by questions like "Who am I? Where do I come from? Where am I going?" (Mitrushchenkova,, 2023). Historically, philosophers like Aristotle have recognized that Identity is shaped both biologically and socially—humans are not merely biological entities but also subjects within social contexts (Mitrushchenkova,, 2023). In contemporary discourse, the concept of Identity has gained significance with developments in psychiatry, psychology, and sociology. For example, Turkle, (1995) in "Life on the Screen" pioneered the exploration of the relationship between technology and identity by investigating how digital technologies interpret and describe human behavior. Since then, extensive research shows that digital technologies not only shape our identities but also influence how we interact with them (Turkle,, 1999; Thorne et al.,, 2015; Adhiarso et al.,, 2019; Hoehe and Thibaut,, 2020; Golubeva,, 2020; Rani,, 2022; Romeo,, 2024). Later,Ching and Foley, (2012) in "Constructing the Self in a Digital World" argues that Identity is a dynamic framework and an continuously evolving, fluid process within digital contexts. Digital technologies such as blogs, social networks, and virtual worlds, not only facilitate the exploration and expression of our identities but also actively shape them. Research focusing on platforms like Facebook illustrates how these tools enable users to overcome physical limitations, thereby enabling the creation of identities that would otherwise be unattainable (Zhao et al.,, 2008; Gardner and Davis,, 2013). Moreover, emerging technologies such as virtual reality (VR), augmented reality (AR), artificial intelligence (AI), and blockchain—key components of metaverse technologies—radically transform our perceptions, experiences, and interactions with the world (Lee et al.,, 2021; Mystakidis,, 2022; Lee et al., 2023b, ; WEF,, 2024). As the distinction between digital and real-world identities becomes increasingly indistinct, understanding the interaction between digital and real-world identities, as well as the changes brought to them by the metaverse, is paramount.
@@
Identity in the metaverse inherits this duality from the real world, yet it becomes even more complex. the metaverse provides users with a creative space unrestricted by physical rules, allowing them to interact with other individuals and environments across various dimensions (Lee et al.,, 2021; Wider et al.,, 2023). This freedom enables users to craft unique digital experiences, unleash creativity, and rethink their achievements, desires, and identities (Mitrushchenkova,, 2023). However, the complexity of identities in the metaverse emerges not only from their multi-layered nature but also from their dynamic evolution and diversity. Users can choose to present different identities in various virtual environments and even continuously adjust and change identity characteristics within the same environment (Mitrushchenkova,, 2023). Such fluidity and plasticity offer users unprecedented freedom of self-expression but also pose significant privacy and security challenges (Ball,, 2020; WEF,, 2024). Neglecting identity consideration in the metaverse can restrict the range of social interactions and degrade the overall user experience, as it limits the ability to represent individuals diversely and safely (Hennig-Thurau et al.,, 2022; Al-kfairy et al.,, 2024; Bradley and Muñoz,, 2024). Furthermore, from a policy and governance perspective, overlooking identity nuances or novel characteristics in the metaverse risks reinforcing existing hegemonic norms and anthropocentric biases, thereby perpetuating real-world biases into the metaverse. This not only impacts technology design and its intended users but also influences the societal norms and values that these technologies reinforce and propagate (Al-kfairy et al.,, 2024; WEF,, 2024).
@@
Metaverse Identity, akin to the broader concept of Identity, is a multi-layered and multi-dimensional construct that has been examined from various disciplinary and perspectives in the literature. It is predominantly considered a foundational term that underpins discussions across diverse topics in specific disciplines, such as philosophy (Spence,, 2008; Van der Merwe,, 2021; Barbara and Haahr,, 2022; Mitrushchenkova,, 2023), psychology (Barbara and Haahr,, 2022; Kim et al.,, 2023; Yang et al., 2024b, ), technology (Falchuk et al.,, 2018; Yang et al.,, 2023; Wang et al.,, 2024; Zichichi et al.,, 2023; Tao et al.,, 2023), management (Banaeian Far and Hosseini Bamakan,, 2023; Wang and Wang,, 2023; Truong et al.,, 2023; Bao et al.,, 2024), and law (Cheong,, 2022; Mitrushchenkova,, 2023; Wu and Zhang,, 2023), which address issues and propose targeted solutions. Current research on Metaverse Identity can also be analyzed from multiple perspectives. The representation perspective treats it as the user’s avatar in the metaverse, focusing on visual representation and interaction (Spence,, 2008; Lee et al.,, 2021; Park and Kim, 2022b, ; Saker and Frith,, 2022; Šímová et al.,, 2023; Lee et al., 2023b, ; Lee et al., 2023a, ). The data perspective considers it as the user’s digital footprint and data aggregation, emphasizing data properties and the extraction of value, including specific identifiers, credentials, and biometric elements (Falchuk et al.,, 2018; Yang et al.,, 2023; Zichichi et al.,, 2023; Tao et al.,, 2023). The social perspective views it through the lens of the user’s social roles and relationships within the metaverse, highlighting social attributes and their interactive effects (Calongne et al.,, 2013; Hollensen et al.,, 2022; Šímová et al.,, 2023; Yang et al., 2024b, ). The economic perspectivedefines it as the user’s economic entity and value carrier, often linked to blockchain technology, with a focus on asset properties and transactional potential (Yang et al.,, 2022; Momtaz,, 2022; Lee et al., 2023b, ; Truong et al.,, 2023; Tao et al.,, 2023; Wang et al.,, 2024). Although these perspectives are not entirely independent and exhibit some overlap, they collectively address the primary facets of the discourse. Some researchers adopt an integrative approach, viewing Metaverse Identity as an extension of real-world identity into the virtual realm, thus emphasizing its multifaceted nature (Saker and Frith,, 2022; Barbara and Haahr,, 2022; Won and Davis,, 2023). This array of perspectives, while not exhaustive, forms the core of the scholarly discourse on this subject. However, research specifically focused on the structured framework and principle of Metaverse Identity on its intrinsic characteristics and impacts remains limited. Establishing such a framework and principle is strategically significant, as recognized by WEF, because it facilitates consensus among a broad range of stakeholders and provides guidance for responsible development in the metaverse.
@@
Under this context, WEF emphasize that Metaverse Identity is an extension of identity as it is known today, encompassing representation, data, and identification (ID). Representation in the metaverse transcends traditional static profile pictures to include customizable digital assets such as avatars, AR fliters, and accessories. These elements reflects various identity aspects from cultural attire to abstract designs, blending physical and digital worlds across 2D screens, AR, and MR. Data points, enhanced by artificial intelligence/machine learning models, are capable of describing and generating identity. They achieve this by analyzing user interactions, movements, and preferences, capturing current activities, predicting future actions, and shaping perceptions. ID systems have evolved from traditional methods like passports and driver’s licenses to unique avatar designs, body-based attestations, or virtual signatures. These innovations validate existence and grant access to specific realms or activities within the metaverse. WEF also note that these three components extends beyond a tangible human to include other digital entities such as chatbots, avatars, and digital replicas. These entitties exhibit varying levels of interaction, autonomy, and behavior within digital experiences, playing vital roles in enabling and enhancing digital interactions (WEF,, 2024). While this framework significantly deepens our understanding of the structural elements of Metaverse Identity, it still lacks the principled guidance necessary for effectively understanding intrinsic characteristics as well as its broader impacts and challenges. This research aims to bridge this gap.

### v2

# Digital Self Governance: Principles and Challenges for Metaverse Identity

@@ Abstract
This paper explores the core principles that should guide the construction and governance of identity in the metaverse and identifies the critical challenges that need to be addressed. Drawing on multidisciplinary theories and perspectives, we define metaverse identity and propose two core principles for understanding its intrinsic characteristics and impacts: Equivalence and Alignment, and Fusion and Expansiveness. The first principle asserts that metaverse identities should align with real-world norms and standards, which is crucial for establishing guidelines and safeguarding rights. The second principle emphasizes the need for seamless integration and boundless expansion of metaverse identities, transcending real-world limitations to accommodate diverse needs and foster inclusive participation. We argue that these two principles are vital for ensuring the accountability, inclusiveness, and consistency in the emerging metaverse era. Additionally, we identify five critical challenges: Identity Interoperability, Legal Implications, Privacy and Identity Management, Deepfakes and Synthetic Identities, and Identity Fragmentation and Psychological Well-being. We discuss potential strategies to navigate these challenges. The paper concludes by underscoring the importance of a proactive and collaborative approach to shaping the future of metaverse identity. As the metaverse continues to evolve, it is imperative that we understand and address the principles and challenges surrounding identity in this uncharted territory and work collectively to build a metaverse that fosters responsible identity construction and expression.
@@
Metaverse, Artificial Intelligent, Identity, Governance, Principles, Interoperability, Privacy, Deepfakes, Synthetic Identity, Identity Fragmentation
journal: Telecommunications Policy
@@
This study articulates two fundamental principles for the governance of identity in the metaverse: “Equivalence and Alignment” and “Fusion and Expansiveness.”
Furthermore, the paper systematically identifies and examines five critical challenges closely related to identity issues in the metaverse.
@@ 1 Introduction
The metaverse is envisioned as the future of the Internet, aspiring to offer a spatial and social Internet experience that seamlessly blend the physical and digital realms using both existing and emerging technologies (Venugopal et al.,, 2023; Hollensen et al.,, 2022; Koohang, A. and Nord, J. and Ooi, K. and Tan, G. and Al-Emran, M. and Aw, E. and Baabdullah, A. and Buhalis, D. and Cham, T. and Dennis, C. and Dutot, V. and Dwivedi, Y. and Hughes, L. and Mogaji, E. and Pandey, N. and Phau, I. and Raman, R. and Sharma, A. and , M. and Ueno, A. and Wong, L.,, 2023). Over the past few decades, and particularly in the last three years, scholarly research on the metaverse has extensively covered various dimensions, including technological advancements (Park and Kim, 2022b, ; Xu et al.,, 2022; Huynh-The et al.,, 2022; Fu et al.,, 2023), ecosystem dynamics (Lee et al.,, 2021), social impacts (Hennig‐Thurau et al.,, 2022; Park and Kim, 2022b, ; Šímová et al.,, 2023), economic potential (Bhattacharya et al.,, 2023), and user behavior and experiences (Riva and Wiederhold,, 2022; Park and Kim, 2022a, ; Buhalis et al.,, 2022).
@@
However, the development and adoption of the metaverse also presents new challenges. In January, the International Criminal Police Organization(INTERPOL) released a research report titled “Metaverse - A Law Enforcement Perspective.” This report underscores that, despite its nascent stages, the metaverse already poses urgent technological, social, and ethical challenges that demands immediate attention. Notably, the metaverse has facilitated the emergence “meta-crimes,” new types of criminal activities that are increasingly concerning as immersive worlds become more integral to our daily lives. Moreover, the report highlight significant governance and policy issues, with the construction of identity within the metaverse being a the primary one. It notes that the anonymity of identities within the metaverse has been exploited for illegal activities such as trafficking stolen goods or promoting unlawful behavior. Additionally, the legal status of avatars remains undefined, prompting the international community to actively explore their legal responsibilities and definitions (INTERPOL,, 2024).
@@
In March, the World Economic Forum(WEF) released a report entitled “Metaverse Identity: Defining the Self in a Blended Reality.” This report elucidates that the metaverse acts as a conduit merging the digital and physical worlds, thereby revolutionizing interactions with information, individuals, and environments. Within this transformative context, identity stands out as a critical element, essential for promoting inclusivity, fairness, accessibility, security, and privacy within the metaverse. The report warns that neglecting the evolution of identity in the metaverse could result in the perpetuation of existing internet flaws. Poorly designed metaverse identities could adversely affect social mobility in the physical world—amidst our increasing reliance on digital platforms—and hinder the secure and private identification, authentication, and verification of individuals or digital entities. Furthermore, these identities may unintentionally endow technology with human-like attributes, cultivating excessive trust and potentially leading to various psychological and emotional harms (WEF,, 2024).
@@
This paper aims to systematically review academic research concerning Metaverse Identity, to explore its intrinsic characteristics and impacts, and to propose core principles that aid stakeholders in intuitively understanding both Metaverse Identity and its broader implications. Additionally, this study seeks to identify the current critical challenges, aiming to provide theoretical references and insights for technological innovations, ethical standards, and policy developments related to Metaverse Identity. Employing a comprehensive, primarily qualitative methodology, this study will integrate literature reviews, textual analysis, comparative studies, and case studies to construct a multidimensional analytical framework. This diversified methodological approach, proven effective for complex subjects, aligns with strategies used by other scholars in studies of virtual identities (Rani,, 2022; Mitrushchenkova,, 2023). Acknowledging the inherent limitations due to the expansive scope of our analysis, this paper seeks to catalyze further interdisciplinary dialogue and research. It is envisioned that future studies will build on this groundwork to provide more profound analyses and robust solutions.
@@ 2 Literature Review
The investigation of Identity represents a crucial philosophical inquiry into human existence. At a fundamental level, Identity encompasses personal details about who they are such as name, age, gender, place of origin, and nationality. However, the Cambridge Dictionary expands this definition to include it as “the fact of being, or feeling that you are, a particular type of person, organization, etc.; the qualities that make a person, organization, etc. different from others” [^1] https://dictionary.cambridge.org/dictionary/english/identity. This broader perspective aligns with philosophical explorations of the ‘Self’, epitomized by questions like “Who am I? Where do I come from? Where am I going?” (Mitrushchenkova,, 2023). Historically, philosophers like Aristotle have recognized that Identity is shaped both biologically and socially—humans are not merely biological entities but also subjects within social contexts (Mitrushchenkova,, 2023). In contemporary discourse, the concept of Identity has gained significance with developments in psychiatry, psychology, and sociology. For example,  Turkle, (1995) in “Life on the Screen” pioneered the exploration of the relationship between technology and identity by investigating how digital technologies interpret and describe human behavior. Since then, extensive research shows that digital technologies not only shape our identities but also influence how we interact with them (Turkle,, 1999; Thorne et al.,, 2015; Adhiarso et al.,, 2019; Hoehe and Thibaut,, 2020; Golubeva,, 2020; Rani,, 2022; Romeo,, 2024). Later, Ching and Foley, (2012) in “Constructing the Self in a Digital World” argues that Identity is a dynamic framework and an continuously evolving, fluid process within digital contexts. Digital technologies such as blogs, social networks, and virtual worlds, not only facilitate the exploration and expression of our identities but also actively shape them. Research focusing on platforms like Facebook illustrates how these tools enable users to overcome physical limitations, thereby enabling the creation of identities that would otherwise be unattainable (Zhao et al.,, 2008; Gardner and Davis,, 2013). Moreover, emerging technologies such as virtual reality (VR), augmented reality (AR), artificial intelligence (AI), and blockchain—key components of metaverse technologies—radically transform our perceptions, experiences, and interactions with the world (Lee et al.,, 2021; Mystakidis,, 2022; Lee et al., 2023b, ; WEF,, 2024). As the distinction between digital and real-world identities becomes increasingly indistinct, understanding the interaction between digital and real-world identities, as well as the changes brought to them by the metaverse, is paramount.
@@
Identity in the metaverse inherits this duality from the real world, yet it becomes even more complex. The metaverse provides users with a creative space unrestricted by physical rules, allowing them to interact with other individuals and environments across various dimensions (Lee et al.,, 2021; Wider et al.,, 2023). This freedom enables users to craft unique digital experiences, unleash creativity, and rethink their achievements, desires, and identities (Mitrushchenkova,, 2023). However, the complexity of identities in the metaverse emerges not only from their multi-layered nature but also from their dynamic evolution and diversity. Users can choose to present different identities in various virtual environments and even continuously adjust and change identity characteristics within the same environment (Mitrushchenkova,, 2023). Such fluidity and plasticity offer users unprecedented freedom of self-expression but also pose significant privacy and security challenges (Ball,, 2020; WEF,, 2024). Neglecting identity consideration in the metaverse can restrict the range of social interactions and degrade the overall user experience, as it limits the ability to represent individuals diversely and safely (Hennig‐Thurau et al.,, 2022; Al-kfairy et al.,, 2024; Bradley and Muñoz,, 2024). Furthermore, from a policy and governance perspective, overlooking identity nuances or novel characteristics in the metaverse risks reinforcing existing hegemonic norms and anthropocentric biases, thereby perpetuating real-world biases into the metaverse. This not only impacts technology design and its intended users but also influences the societal norms and values that these technologies reinforce and propagate (Al-kfairy et al.,, 2024; WEF,, 2024).
Metaverse Identity, akin to the broader concept of Identity, is a multi-layered and multi-dimensional construct that has been examined from various disciplinary and perspectives in the literature. It is predominantly considered a foundational term that underpins discussions across diverse topics in specific disciplines, such as philosophy (Spence,, 2008; Van der Merwe,, 2021; Barbara and Haahr,, 2022; Mitrushchenkova,, 2023), psychology (Barbara and Haahr,, 2022; Kim et al.,, 2023; Yang et al., 2024b, ), technology (Falchuk et al.,, 2018; Yang et al.,, 2023; Wang et al.,, 2024; Zichichi et al.,, 2023; Tao et al.,, 2023), management (Banaeian Far and Hosseini Bamakan,, 2023; Wang and Wang,, 2023; Truong et al.,, 2023; Bao et al.,, 2024), and law (Cheong,, 2022; Mitrushchenkova,, 2023; Wu and Zhang,, 2023), which address issues and propose targeted solutions.
@@
Current research on Metaverse Identity can also be analyzed from multiple perspectives. The representation perspective treats it as the user’s avatar in the metaverse, focusing on visual representation and interaction (Spence,, 2008; Lee et al.,, 2021; Park and Kim, 2022b, ; Saker and Frith,, 2022; Šímová et al.,, 2023; Lee et al., 2023b, ; Lee et al., 2023a, ). The data perspective considers it as the user’s digital footprint and data aggregation, emphasizing data properties and the extraction of value, including specific identifiers, credentials, and biometric elements (Falchuk et al.,, 2018; Yang et al.,, 2023; Zichichi et al.,, 2023; Tao et al.,, 2023). The social perspective views it through the lens of the user’s social roles and relationships within the metaverse, highlighting social attributes and their interactive effects (Calongne et al.,, 2013; Hollensen et al.,, 2022; Šímová et al.,, 2023; Yang et al., 2024b, ). The economic perspective defines it as the user’s economic entity and value carrier, often linked to blockchain technology, with a focus on asset properties and transactional potential (Yang et al.,, 2022; Momtaz,, 2022; Lee et al., 2023b, ; Truong et al.,, 2023; Tao et al.,, 2023; Wang et al.,, 2024). Although these perspectives are not entirely independent and exhibit some overlap, they collectively address the primary facets of the discourse. Some researchers adopt an integrative approach, viewing Metaverse Identity as an extension of real-world identity into the virtual realm, thus emphasizing its multifaceted nature (Saker and Frith,, 2022; Barbara and Haahr,, 2022; Won and Davis,, 2023). This array of perspectives, while not exhaustive, forms the core of the scholarly discourse on this subject.
Within these contexts, Metaverse Identity can be defined as the multi-dimensional construct of a user’s digital self, integrating personal representative attributes, data footprints, social roles, and economic elements within the metaverse. This construct encompasses various facets of an individual’s presence and activities within the metaverse, and encapsulates the complex interplay between an individual’s online persona and their real-world identity, reflecting a broad spectrum of personal and social facets. WEF also note that Metaverse Identity extends beyond a tangible human to include other digital entities such as chatbots, avatars, and digital replicas. These entitties exhibit varying levels of interaction, autonomy, and behavior within digital experiences, playing vital roles in enabling and enhancing digital interactions (WEF,, 2024).
@@
However, research specifically focused on the structured framework and principles of Metaverse Identity, particularly regarding its intrinsic characteristics and impacts, remains limited. Establishing such a framework is strategically significant because it facilitates consensus among a broad range of stakeholders and provides guidance for responsible development in the metaverse (WEF,, 2024). WEF proposed its framework that encompass representation, data, and identification (ID). Representation in the metaverse transcends traditional static profile pictures to include customizable digital assets such as avatars, AR fliters, and accessories. These elements reflects various identity aspects from cultural attire to abstract designs, blending physical and digital worlds across 2D screens, AR, and MR. Data points, enhanced by artificial intelligence/machine learning models, are capable of describing and generating identity. They achieve this by analyzing user interactions, movements, and preferences, capturing current activities, predicting future actions, and shaping perceptions. ID systems have evolved from traditional methods like passports and driver’s licenses to unique avatar designs, body-based attestations, or virtual signatures. These innovations validate existence and grant access to specific realms or activities within the metaverse. While this framework enhances our understanding of Metaverse Identity, it still lacks the principled guidance necessary for effectively understanding intrinsic characteristics as well as its broader impacts and challenges. This research aims to bridge this gap.